# A Generalized Analytical Framework for the Nonlinear Best-Worst Method

### v1

@@ Abstract
To eliminate the need for optimization software in calculating weights using the nonlinear model of the Best-Worst Method (BWM), Wu et al. proposed an analytical framework for deriving optimal interval-weights. They also introduced a secondary objective function to select the best optimal weight set. However, their framework is only compatible with a single Decision-Maker (DM) and preferences quantified using the Saaty scale. In this research, we generalize their framework to accommodate any number of DMs and any scale. We first derive an analytical expression for optimal interval-weights and select the best optimal weight set. After demonstrating that the values of Consistency Index (CI) for the Saaty scale in the existing literature are not well-defined, we derive n formula for computing CI. We also derive analytical expressions for the Consistency Ratio (CR), enabling its use as an input-based consistency indicator and proving that CR satisfies some key properties, ensuring its reliability as a consistency indicator. Furthermore, we observe that criteria with equal preferences may get different weights when multiple best/worst criteria are present. To address this limitation, we modify the original optimization model for weight computation in such instances, solve it analytically to obtain optimal interval-weights, and select the best optimal weight set. Finally, we demonstrate and validate the proposed approach using numerical examples.
@@ 1 Introduction
Multi-Criteria Decision-Making (MCDM), a fundamental branch of operations research, addresses complex decision scenarios involving conflicting criteria. Typically, MCDM approaches solve decision problems through two sequential steps: determining weights of decision criteria and ranking of alternatives. This classification divides MCDM methods into two categories: weight determination methods (such as AHP [37], ANP [38] and SMARTS [15]) and ranking methods (including TOPSIS [20], VIKOR [29], ELECTRE [36] and PROMETHEE [5]).
@@
The Best-Worst Method (BWM) is a widely used weight calculation method that utilizes pairwise comparisons between decision criteria to derive weights [32]. In this method, the optimal weights are obtained by minimizing the distance between weight ratios and given comparison values. Based on different distance functions, various models of BWM have been developed. The original model of BWM, proposed by Rezaei [32], uses maximum deviation as the distance function. Since this approach involves solving a nonlinear optimization problem, it is known as the nonlinear BWM. Kocak et al. [22] introduced a model of BWM based on Euclidean distance, termed the Euclidean BWM. Brunelli and Rezaei [8] proposed the multiplicative BWM, incorporating the metric $\max\{\nicefrac{{x}}{{y}},\nicefrac{{y}}{{x}}\}$ defined on $(\mathbb{R_{+}},\cdot,\leq)$ into the BWM framework. Amiri and Emamat [3] further developed a goal programming-based BWM using the taxicab distance. Tu et al. [42] introduced two alternative BWM formulations: the approximate eigenvalue model and the logarithmic least squares model. For group decision-making, Safarzadeh et al. [39] proposed two extensions, one based on total deviation and the other based on maximum deviation. Xu and Wang [48] developed eleven models for individual DMs and nine for group decision-making contexts. Mohammadi and Rezaei [28] incorporated Bayesian into the BWM framework, establishing a probabilistic approach to group decision-making. Corrente et al. [11] developed the parsimonious BWM, an enhanced version of the nonlinear BWM specifically designed for decision contexts involving numerous alternatives.
@@
To address the non-uniqueness of optimal weight sets in the nonlinear BWM, Rezaei [33] derived optimal interval-weights by formulating two optimization models. Later, he introduced the concentration ratio to measure the dispersion of these interval-weigths [34]. He also developed the linear BWM [33], which retains the underlying philosophy of the nonlinear BWM while transforming its optimization framework into a linear formulation. While this model guarantees a unique weight set, its feasible region differs from the original nonlinear approach. Wu et al. [47] developed an analytical approach to derive optimal interval-weights for the nonlinear BWM, eliminating the model’s dependency on optimization software. Building on this foundation, they incorporated a secondary objective function to determine the best optimal weight set from the solution space. Ratandhara and Kumar [31] subsequently proposed an analytical framework for the multiplicative BWM, achieving the same objectives of software independence and selection of the best optimal weight set.
@@
Consistency measurement of decision data is crucial in MCDM methods since outcomes depend directly on this input. In BWM, consistency evaluation is performed through the Consistency Ratio (CR), which is typically computed using the optimal objective value and Consistency Index (CI) [32]. This output-based consistency indicator can only provide feedback about inconsistencies after completing all calculations, resulting in reduced time efficiency. Liang et al. [26] introduced an alternative input-based consistency indicator called input-based CR, along with establishing threshold values for both output-based and input-based CR to check admissibility of preference values. Furthermore, Lei et al. [24] developed an optimization model that recommends optimal preference modifications while achieving both ordinal consistency and an acceptable level of cardinal consistency.
@@
To handle uncertain preferences, several fuzzy-set-based extensions of BWM have been proposed. Guo and Zhao [17] extended the nonlinear BWM to a fuzzy environment, while Rostami et al. [35] introduced a fuzzy adaptation of the goal programming-based BWM. Additionally, Ratandhara and Kumar [30] proposed an $\alpha$-cut interval-based model of fuzzy BWM. The BWM framework has also been extended to more advanced uncertainty representations, including intuitionistic fuzzy sets [46, 45, 10], hesitant fuzzy sets [2, 25] and spherical fuzzy sets [19]. Moreover, BWM has been integrated with other MCDM techniques, such as BWM-VIKOR [1, 16], BWM-ELECTRE [49, 9], BWM-TOPSIS [44, 41], BWM-MULTIMOORA [18, 50] and Best-Worst Tradeoff (BWT) method [27]. Owing to its simplicity and reliability, BWM has been widely applied in practical decision-making problems, notably in location selection [43, 4], logistics risk assessment [12, 13] and supplier selection [40, 44].
@@
    The framework proposed by Wu et al. [47] works well when preferences are quantified using the Saaty scale, but leads to different optimal interval-weights and a best optimal weight set than their actual values (in fact, optimal interval-weights are not well-defined) when some other scale such as the Salo-Hämäläinen, Lootsma or Donegan-Dodd-McMaster scale is used.
@@
    The existing framework leads to non-well-defined optimal interval-weights and a different best weight set than the actual one, even when preferences are quantified using the Saaty scale in the presence of multiple DMs.
@@
    In the presence of multiple best/worst criteria, computing weights by arbitrarily selecting any one best/worst criterion leads to different weights (both interval-weights and the best weights) for criteria having equal preference.
In this work, we propose a generalized analytical framework for the nonlinear BWM that is compatible with any preference scale and any number of DMs. Our approach derives optimal interval-weights by solving an optimal modification-based optimization problem, which is analytically equivalent to the original optimization model. From the collection of all optimal weight sets, we also select the best optimal weight set. Based on this framework, we derive an exact formula for the CI valid for any scale, along with an analytical expression for the CR that serves as an input-based consistency indicator. We further establish essential properties of CR to verify its validity as a consistency measure. To ensure equal weights for criteria with equal preferences when multiple best/worst criteria exist, we develop a modified optimization model and derive corresponding optimal interval-weights and the best optimal weight set analytically. We validate the proposed approach and demonstrate its effectiveness through numerical examples.
The remainder of this paper is organized as follows. Section 2 introduces fundamental definitions and provides a concise overview of the nonlinear BWM and its existing framework. Section 3 identifies and examines critical research gaps in the current methodology. In Section 4, we present our generalized analytical framework for the nonlinear BWM, supported by demonstrative numerical examples. Finally, Section 5 discusses concluding remarks and suggests potential directions for future research.
2 Preliminary
In this section, we discuss some fundamental definitions and key results, along with a brief overview of the nonlinear BWM and its analytical framework.
@@ 2.1 Basic Concepts and Results
Let $C=\{c_{1},c_{2},\ldots,c_{n}\}$ be the set of decision criteria, and let $D=\{c_{1},c_{2},\ldots,c_{n}\}\setminus\{c_{b},c_{w}\}$ throughout the article. Whenever unambiguous, we use the abbreviated notations $C=\{1,2,\ldots,n\}$ and $D=\{1,2,\ldots,n\}\setminus\{b,w\}$.
The Pairwise Comparison System (PCS) is the pair $(A_{b},A_{w})$, where $A_{b}=(a_{b1},a_{b2},\ldots,a_{bn})$ is the best-to-other vector and $A_{w}=(a_{1w},a_{2w},\ldots,a_{nw})^{T}$ is the other-to-worst vector. Here, $a_{ij}$ represents the relative preference of the $i^{th}$ criterion over the $j^{th}$ criterion.
@@ 2.2 Nonlinear BWM
The BWM is an MCDM technique that derives criteria weights through pairwise comparisons between the best (most preferable), the worst (least preferable) and other criteria [32]. The steps of the BWM are as follows.
Step 1: Formation of the set of decision criteria $C=\{c_{1},c_{2},\ldots,c_{n}\}$.
Furthermore, We adopt the notation $D=\{c_{1},c_{2},\ldots,c_{n}\}\setminus\{c_{b},c_{w}\}$ throughout this work. When no ambiguity arises, we simplify the notation to $C=\{1,2,\ldots,n\}$ and $D=\{1,2,\ldots,n\}\setminus\{b,w\}.$
@@
Step 3: Determination of the best-to-other vector $A_{b}=(a_{b1},a_{b2},\ldots,a_{bn})$ and the other-to-worst vector $A_{w}=(a_{1w},a_{2w},\ldots,a_{nw})^{T}$.
The preferences $a_{ij}$ are typically expressed as linguistic terms, which are then quantified using established scales as shown in Table 1.
@@
The effectiveness of an MCDM method depends on the decision data that is often inconsistent because of human engagement. A key requirement for any rigorous MCDM methodology is the ability to assess and quantify these inconsistencies. In the BWM, this assessment is performed using the Consistency Ratio (CR) defined as
@@ 2.3 An Analytical Framework for the Nonlinear BWM
Wu et al. [47] proposed an analytical approach to derive optimal interval weights without requiring optimization software. They also introduced a secondary objective function to select the best optimal weight set from the collection of all optimal weight sets.
2.3.1 Calculation of Interval-Weights
Consider the following optimization model, driven by the optimal modification of PCS.
@@ 2.3.1 Calculation of Interval-Weights
This implies that problem (7) is an equivalent to problem (3). Consequently, the analytical solution of problem (3) can be obtained by solving problem (7) analytically. To describe the analytical solution of Problem (7), some mathematical symbols must first be defined.
Let $D_{1}=\{i\in D:a_{bi}\times a_{iw}<a_{bw}\}$, $D_{2}=\{i\in D:a_{bi}\times a_{iw}>a_{bw}\}$ and $D_{3}=\{i\in D:a_{bi}\times a_{iw}=a_{bw}\}$.
@@
    $(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i})=a_{bw}-\epsilon_{i}$ (10)
@@
Let $i_{1}\in D_{1}$ and $j_{0}\in D_{2}$ be such that $\epsilon_{i_{1}}=\max\{\epsilon_{i}:i\in D_{1}\}$ and $\epsilon_{i_{2}}=\max\{\epsilon_{i}:i\in D_{2}\}$ respectively. Then, by [47, Proposition 3], we get
@@
Now, by [47, Theorem 3], the collection of all optimally modified PCS is
@@
Now, from [47, Theorem 4], the collection of all optimal weights of criterion $c_{i}$ is ${w_{i}^{l}}^{*},{w_{i}^{u}}^{*}]$, where
${w_{i}^{l}}^{*}=\frac{\inf\{\tilde{a}_{iw}^{*}\}}{\inf\{\tilde{a}_{iw}^{*}\}+\displaystyle\sum_{\begin{subarray}{c}j\in C\\ j\neq i\end{subarray}}\sup\{\tilde{a}_{jw}^{*}\}},\quad{w_{i}^{u}}^{*}=\frac{\sup\{\tilde{a}_{iw}^{*}\}}{\sup\{\tilde{a}_{iw}^{*}\}+\displaystyle\sum_{\begin{subarray}{c}j\in C\\ j\neq i\end{subarray}}\inf\{\tilde{a}_{jw}^{*}\}},$ (19)
where $i\in C$, and thus,
@@
2.3.2 A Secondary Objective Function to Obtain a Unique Weight Set
Since the nonlinear BWM may yield multiple optimal weight sets—and while such multiplicity is desirable in some cases—DMs usually prefer a unique weight set. To address this, Wu et al. [47] introduced a secondary objective function to obtain the best optimal weight set.
In this approach, an optimally modified PCS having the minimum value of $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}$ for all $i\in C$ is first selected as the best optimally modified PCS. As shown in [47, Theorem 5], for $i\in D$, the minimum possible value of $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}$ is the optimal objective value of minimization problem

### v2

@@ Abstract
The nonlinear model of the best-worst method frequently produces multiple optimal weight sets, which are conventionally determined through optimization software. While an analytical approach exists that provides both a closed-form expression for the optimal interval-weights and a secondary objective function to determine the best optimal weight set, we demonstrate that this approach is only valid when preferences are quantified using the Saaty scale and only a single decision-maker is involved. To tackle this issue, we propose a framework compatible with any scale and any number of decision-makers. We first derive an analytical expression for optimal interval-weights and then select the best optimal weight set. After demonstrating that the values of consistency index for the Saaty scale in the existing literature are not well-defined, we derive a formula of consistency index. We also obtain an analytical expression for the consistency ratio, enabling its use as an input-based consistency indicator. Furthermore, we establish that when multiple best/worst criteria are present, weights may vary among best criteria and among the worst criteria. To address this limitation, we modify the original optimization model for weight computation in such instances, solve it analytically to obtain optimal interval-weights and then select the best optimal weight set using a secondary objective function. Finally, we demonstrate and validate the proposed approach using numerical examples and a real-world case study of ranking barriers to energy efficiency in buildings.
@@ 1 Introduction
Multi-Criteria Decision-Making (MCDM), a fundamental branch of operations research, addresses complex decision scenarios involving conflicting criteria. Typically, MCDM approaches solve decision problems through two sequential steps: determining weights of decision criteria and ranking of alternatives. This classification divides MCDM methods into two categories: weight determination methods (such as AHP [48], ANP [49], and SMARTS [18]) and ranking methods (including TOPSIS [26], VIKOR [38], ELECTRE [46], and PROMETHEE [5]).
@@
The Best-Worst Method (BWM) is a widely used weight calculation method that utilizes pairwise comparisons between decision criteria to derive weights [42]. In this method, the optimal weights are obtained by minimizing the distance between weight ratios and given comparison values. Based on different distance functions, various models of BWM have been developed. The original model of BWM, proposed by Rezaei [42], uses maximum deviation as the distance function. Since this approach involves solving a nonlinear optimization problem, it is known as the nonlinear BWM. Kocak et al. [28] introduced a model of BWM based on Euclidean distance, termed the Euclidean BWM. Brunelli and Rezaei [8] proposed the multiplicative BWM, incorporating the metric $\max\{\nicefrac{{x}}{{y}},\nicefrac{{y}}{{x}}\}$ defined on $(\mathbb{R_{+}},\cdot,\leq)$ into the BWM framework. Amiri and Emamat [3] developed a goal programming-based BWM using the taxicab distance. Tu et al. [53] introduced two alternative BWM formulations: the approximate eigenvalue model and the logarithmic least squares model. For group decision-making, Safarzadeh et al. [50] proposed two models of BWM, one based on total deviation and the other based on maximum deviation. Xu and Wang [60] discussed eleven BWM models for individual Decision-Makers (DMs) and nine for group decision-making contexts. Mohammadi and Rezaei [37] incorporated Bayesian into the BWM framework, establishing a probabilistic approach to group decision-making. Corrente et al. [14] developed the parsimonious BWM, an enhanced version of the nonlinear BWM specifically designed for decision contexts involving numerous alternatives.
@@
To address the non-uniqueness of optimal weight sets in the nonlinear BWM, Rezaei [43] derived optimal interval-weights using two optimization models. Later, he introduced the concentration ratio to measure the dispersion of these interval-weigths [44]. He also developed the linear BWM [43], which retains the underlying philosophy of the nonlinear BWM while transforming its optimization framework into a linear formulation. While this model guarantees a unique weight set, its feasible region differs from the original nonlinear approach. Wu et al. [59] developed an analytical approach to derive optimal interval-weights for the nonlinear BWM, eliminating the model’s dependency on optimization software. Building on this foundation, they introduced a secondary objective function to determine the best optimal weight set from the solution space. Ratandhara and Kumar [41] subsequently proposed an analytical framework for the multiplicative BWM, achieving the same objectives of software independence and selection of the best optimal weight set.
@@
Measurement of consistency of decision data is crucial in MCDM methods since outcomes depend directly on this input data. In BWM, consistency evaluation is performed through the Consistency Ratio (CR), which is computed using the optimal objective value of optimization model and Consistency Index (CI) [42]. This output-based consistency indicator can only provide feedback about inconsistencies after completing all calculations, resulting in reduced time efficiency. Liang et al. [34] introduced an alternative input-based consistency indicator called input-based CR, along with establishing threshold values for both output-based and input-based CR to check admissibility of preference values. Furthermore, Lei et al. [32] developed an optimization model that recommends optimal preference modifications while achieving both ordinal consistency and an acceptable level of cardinal consistency.
@@
To handle uncertain preferences, several fuzzy-set-based extensions of BWM have been proposed. Guo and Zhao [20] extended the nonlinear BWM to a fuzzy environment, while Rostami et al. [45] introduced a fuzzy adaptation of the goal programming-based BWM. Additionally, Ratandhara and Kumar [40] proposed an $\alpha$-cut interval-based model of fuzzy BWM. The BWM framework has also been extended to more advanced uncertainty representations, including intuitionistic fuzzy sets [57, 56, 13], hesitant fuzzy sets [2, 33] and spherical fuzzy sets [24]. Moreover, BWM has been integrated with other MCDM techniques, such as BWM-VIKOR [1, 19], BWM-ELECTRE [61, 12], BWM-TOPSIS [55, 52], BWM-MULTIMOORA [22, 62] and Best-Worst Tradeoff (BWT) method [35]. Owing to its simplicity and reliability, BWM has been widely applied in practical decision-making problems, notably in location selection [54, 4], logistics risk assessment [15, 16], and supplier selection [51, 55].
@@
    The framework proposed by Wu et al. [59] works well when preferences are quantified using the Saaty scale, but leads to different optimal interval-weights and a best optimal weight set than their actual values (in fact, non-well-defined optimal interval-weights) when some other scale such as the Salo-Hämäläinen, Lootsma or Donegan-Dodd-McMaster scale is used. This is a significant concern, as the quantification of preferences using the Saaty scale may not always be reliable [29].
@@
    The existing framework leads to non-well-defined optimal interval-weights and a different best weight set than the actual one in the presence of multiple DMs even when preferences are quantified using the Saaty scale.
@@
    In the presence of multiple best/worst criteria, computing weights by arbitrarily selecting any one best/worst criterion leads to weight differences (in both interval-weights and the best weights) among the best criteria and among the worst criteria.
In this work, we propose an analytical framework for the nonlinear BWM that is compatible with any preference scale and any number of DMs. Our approach derives optimal interval-weights by solving an optimal modification-based optimization problem, which is equivalent to the original optimization model. From the collection of all optimal weight sets, we also select the best optimal weight set by introducing a secondary objective function. Based on this framework, we derive a formula for computing CI that is valid for any scale, along with an analytical expression for CR that serves as an input-based consistency indicator. We further establish essential properties of CR to verify its reliability as a consistency measure. To ensure equal weights for all best and all worst criteria in cases with multiple best/worst criteria, we develop a modified optimization model and analytically derive both the corresponding optimal interval-weights and the best optimal weight set. We validate the proposed approach through numerical examples. Additionally, we rank barriers to energy efficiency in buildings to demonstrate applicability of the framework.
The remainder of this paper is organized as follows. Section 2 introduces fundamental definitions and provides a concise overview of the nonlinear BWM and its existing framework. Section 3 identifies and examines critical research gaps in the current methodology. Section 4 presents a generalized analytical framework for the nonlinear BWM including weight calculation, consistency analysis, and numerical examples. Section 5 demonstrates the applicability of the proposed approach by ranking barriers to energy efficiency in buildings. Finally, Section 6 discusses concluding remarks and suggests potential directions for future research.
2 Preliminaries
In this section, we discuss some fundamental definitions and results, along with a brief overview of the nonlinear BWM and its analytical framework.
@@ 2.1 Basic Concepts and Results
Let $C=\{c_{1},c_{2},\ldots,c_{n}\}$ be the set of decision criteria, and let $D=\{c_{1},c_{2},\ldots,c_{n}\}\setminus\{c_{b},c_{w}\}$ throughout the article. For brevity, when the context is clear, we simplify the notation by representing the sets as $C=\{1,2,\ldots,n\}$ and $D=\{1,2,\ldots,n\}\setminus\{b,w\}$.
The Pairwise Comparison System (PCS) is the pair $(A_{b},A_{w})$, where $A_{b}=(a_{b1},a_{b2},\ldots,a_{bn})$ and $A_{w}=(a_{1w},a_{2w},\ldots,a_{nw})^{T}$ denote the best-to-other and the other-to-worst vector respectively. Here, $a_{ij}$ represents the relative preference of the $i^{th}$ criterion over the $j^{th}$ criterion.
@@
Intermediate Value Theorem (IVT).
[47] Let $f$ be a real-valued continuous function on $[a,b]$ for which $f(a)<c<f(b)$. Then there is a $a<x_{0}<b$ such that $f(x_{0})=c$.
@@ 2.2 Nonlinear BWM
BWM is a pairwise comparison-based MCDM method for deriving criteria weights [42]. The steps of BWM are as follows.
Step 1: Formation of the set of decision criteria $C$.
@@
Step 3: Determination of the PCS $(A_{b},A_{w})$.
The preferences $a_{ij}$ are generally expressed as linguistic terms, which are then quantified using an established scale. Some such scales are given in Table 1.
@@
The effectiveness of an MCDM method depends on the decision data that is often inconsistent because of human engagement. A key requirement for any MCDM method is the ability to assess and quantify these inconsistencies. In BWM, this assessment is performed using the Consistency Ratio (CR) defined as
@@ 2.3 An Analytical Framework for the Nonlinear BWM
Wu et al. [59] proposed an analytical approach to derive optimal interval-weights without requiring optimization software. They also introduced a secondary objective function to select the best optimal weight set from the collection of all optimal weight sets.
2.3.1 Interval-Weights
The optimal interval-weights are calculated using the following optimization model, which is driven by an optimal modification of the PCS.
@@ 2.3.1 Interval-Weights
This implies that problem (7) is equivalent to problem (3). Consequently, the analytical expression for optimal solutions of problem (3) can be obtained by solving problem (7) analytically. To describe the analytical solution of problem (7), some mathematical symbols must first be introduced.
Let $D_{1}=\{i\in D:a_{bi}\times a_{iw}<a_{bw}\}$, $D_{2}=\{i\in D:a_{bi}\times a_{iw}>a_{bw}\}$, and $D_{3}=\{i\in D:a_{bi}\times a_{iw}=a_{bw}\}$.
@@
    $(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i})=a_{bw}-\epsilon_{i}.$ (10)
@@
Let $i_{1}\in D_{1}$ and $i_{2}\in D_{2}$ be such that $\epsilon_{i_{1}}=\max\{\epsilon_{i}:i\in D_{1}\}$ and $\epsilon_{i_{2}}=\max\{\epsilon_{i}:i\in D_{2}\}$. Then, by [59, Proposition 3], we get
@@
Now, by [59, Theorem 3], the collection of all optimally modified PCS is
@@
From [59, Theorem 4], the optimal interval-weight of criterion $c_{i}$, $i\in C$, is ${w_{i}^{l}}^{*},{w_{i}^{u}}^{*}]$, where
${w_{i}^{l}}^{*}=\frac{\inf\{\tilde{a}_{iw}^{*}\}}{\inf\{\tilde{a}_{iw}^{*}\}+\displaystyle\sum_{\begin{subarray}{c}j\in C\\ j\neq i\end{subarray}}\sup\{\tilde{a}_{jw}^{*}\}},\quad{w_{i}^{u}}^{*}=\frac{\sup\{\tilde{a}_{iw}^{*}\}}{\sup\{\tilde{a}_{iw}^{*}\}+\displaystyle\sum_{\begin{subarray}{c}j\in C\\ j\neq i\end{subarray}}\inf\{\tilde{a}_{jw}^{*}\}}.$ (19)
Therefore
@@
2.3.2 A Secondary Objective Function
To identify the optimal weight set that best preserves the given data, Wu et al. [59] proposed the incorporation of a secondary objective function.
In this approach, an optimally modified PCS having the minimum value of $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}$ for all $i\in C$ is selected as the best optimally modified PCS. As shown in [59, Theorem 5], for $i\in D$, the minimum possible value of $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}$ is the optimal objective value of problem